EXCITONIC COUPLING AND PHENYL RING TORSION IN DIPHENYLMETHANE

\underline{JAIME A. STEARNS}, ALOKE DAS, TALITHA M. SELBY, AND TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN, 47907; DAVID F. PLUSQUELLIC, National Institute of Standards and Technology, Gaithersburg, MD, 20899.

Diphenylmethane is a simple bichromophore in which two phenyl rings are connected only by a methylene group, giving considerable flexibility in the ring torsion coordinates as well as substantial electronic coupling between the rings. According to DFT calculations, the minimum energy structure has C_{2} symmetry with a ring torsion angle near 60^{o}. There are two such minima connected by a C_{2v} structure 176 cm^{−1} higher in energy. The rotationally-resolved fluorescence excitation spectrum of the S_{0}-S_{1} origin indicated a 70% a-type, 30% c-type transition moment, suggestive of an excitonic state in which the zero-order transition moment is rotated significantly from that of toluene. CIS underestimates the a-type character of the band, while time-dependent DFT overestimates it. The ground state experimental rotational constants are consistent with those predicted by DFT. A Franck-Condon progession in the torsional coordinate was observed in the experimental jet-cooled resonant two-photon ionization and dispersed fluorescence spectra. This progression was well-fit using a harmonic Franck-Condon analysis indicating a 3^{o} change in the ring torsion angle upon excitation to S_{1}. The S_{0}-S_{2} origin was assigned to a vibration 123 cm^{−1} above the S_{0}-S_{1} origin. The dispersed fluorescence spectrum from this transition showed a great deal of activity in low frequency vibrations which were not present in the excitation spectrum, indicating that a vibronic band of the S_{1} state may be resonant with the S_{0}-S_{2} origin. To test the effect of asymmetry on the excitonic coupling, the spectroscopy of 4-methyldiphenylmethane has also been studied. Addition of a methyl group to one chromophore completely localizes the electronic excitation, demonstrated by the fact that the two electronic origins are very near those of toluene and para-xylene. Dispersed fluorescence from the S_{0}-S_{2} origin showed only S_{0}-S_{1} origin-like emission broadened by IVR, indicative of extremely rapid electronic energy transfer.